OpenStudy (anonymous):

1. 2(x + 5) > 8x – 8 2. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

@shaik0124

OpenStudy (shaik0124):

2x+10>8x-8 6x>18 x>3

OpenStudy (shaik0124):

@sammy5592

OpenStudy (anonymous):

thanks 2. now please

OpenStudy (shaik0124):

-2x+6>=5-x-3 -2x+x>=5-3-6 -x>=-4 x>=4

OpenStudy (shaik0124):

@sammy5592

OpenStudy (anonymous):

TY so much

OpenStudy (shaik0124):

okkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkk

OpenStudy (shaik0124):

@sammy5592 mdl

OpenStudy (shaik0124):

@sammy5592